ZEKE SPECTROSCOPY OF GALLIUM-METHYLAMINE COMPLEXES: Ga-NH_{n}(CH_{3})_{3}_{−}_{n} (n = 0-2)

\underline{PARAGRANJITA BHOWMIK}, SHENGGANG LI, JASON F. FULLER, GRETCHEN K. ROTH-SCHOPF, BRADFORD R. SOHNLEIN, XU WANG, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Gallium-amine complexes were produced by reactions of gallium atoms and methylamines seeded in helium or argon carrier gases. The Ga-NH_{n}(CH_{3})_{3}_{−}_{n} complexes have adiabatic ionization potentials of 39330, 38790, and 38081 cm^{−1}for n = 2, 1, and 0, respectively. The ZEKE spectrum of Ga-NH_{2}CH_{3} exhibits 299 and 124 cm^{−1}vibrations in the ionic state and a 93 cm^{−1}vibration in the neutral state. In the case of Ga-NH(CH_{3})_{2}, 864, 331, 206, and 132 cm^{−1}vibrations were measured for the ion, and 177 and 128 cm^{−1}vibrations for the neutral molecule. For Ga-N(CH_{3})_{3}, the spectrum displays ionic vibrations of 785, 462, 188, and 111 cm^{−1}and a neutral vibration of 133 cm^{−1}. Assignments of the observed vibrations will be discussed in comparison with theoretical calculations.